# Sensing and Steering Stereotypes: Extracting and Applying Gender Representation Vectors in LLMs

### v2

@@ 6 Conclusion
Our work studies gender representations in LLMs, specifically through the feminine–masculine spectrum. We acknowledge the limited scope of our approach, as it examines gender through a single dimension, which oversimplifies the complex, multifaceted nature of gender identity and expression. Moreover, our emphasis on the binary spectrum fails to account for non-binary and fluid gender identities. Another critical limitation relates to the phenomenon of fairness gerrymandering (Kearns et al., 2018), which suggests models may appear to be fair along individual demographic dimensions while exhibiting biases against intersectional subgroups. Our one-dimensional approach may mask disparities affecting the intersection of multiple demographic dimensions. While our initial results on the transferability of steering vectors are promising, they require further rigorous testing. Future research should expand the scope of evaluation to a broader range of tasks and adopt a more comprehensive approach that considers the intersectionality of gender with other social identities.
@@ E.2 Varying Disparity Score Threshold
Results shown in both Section 3.4 and Section 4.2 are based on the same score threshold $\delta$ of 0.05. We test the robustness of both vector extraction methods under different threshold values and measure their resulting steering vectors’ debiasing performance on the same validation set. We use eight different values of $\delta$ from 0.01 to 0.3 with increasing increments. Figure 19 shows the range of RMS bias scores after debiasing under different $\delta$ across all eight models. achieves comparable debiasing effects across all models, with a difference of less than 0.05 in bias scores for the same model. MD exhibits the largest discrepancy in bias scores for the Llama-3.1-8B model, with a difference of 0.1. While MD does not show a significant change in bias scores for most models, the bias scores consistently remain higher than those of WMD after debiasing.

### v3

@@ 6 Conclusion
Our work studies gender representations in LLMs, specifically through the feminine–masculine spectrum. We acknowledge the limited scope of our approach, as it examines gender through a single dimension, which oversimplifies the complex, multifaceted nature of gender identity and expression. Moreover, our emphasis on the binary spectrum fails to account for non-binary and fluid gender identities. Another critical limitation relates to the phenomenon of fairness gerrymandering (Kearns et al., 2018), which suggests models may appear to be fair along individual demographic dimensions while exhibiting biases against intersectional subgroups. Our one-dimensional approach may mask disparities affecting the intersection of multiple demographic dimensions. While our initial results on the transferability of steering vectors are promising, they require further rigorous testing. Moreover, our bias mitigation method (Section 4.2) may not be applicable to contexts where model predictions should depend on factual gender information. This may require disentangling the representations of factual gender from undesired gender associations or using a “condition vector” to detect the appropriate contexts, similar to Lee et al. (2025). Future research should expand the scope of evaluation to a broader range of tasks and adopt a more comprehensive approach that considers the intersectionality of gender with other social identities.
@@ E.2 Varying Disparity Score Threshold
Results shown in both Section 3.4 and Section 4.2 are based on the same score threshold $\delta$ of 0.05. We test the robustness of both vector extraction methods under different threshold values and measure the debiasing performance of their resulting steering vectors on the same validation set. We use eight different values of $\delta$ from 0.01 to 0.3 with increasing increments. Figure 20 shows the range of RMS bias scores after debiasing under different $\delta$ across all eight models. WMD achieves comparable debiasing effects across all models, with a difference of less than 0.05 in bias scores for the same model. MD exhibits the largest discrepancy in bias scores for the Llama-3.1-8B model, with a difference of 0.1. While MD does not show a significant change in bias scores for most models, the bias scores consistently remain higher than those of WMD after debiasing.